OpenStudy (anonymous):

What is the value of the expression 9 − (1/2 to the 4th power.) ⋅ 48? 1 3 6 9

OpenStudy (anonymous):

the answer is 6

OpenStudy (anonymous):

is this greed peeps?

OpenStudy (anonymous):

\[9-(\frac{ 1 }{ 2})^{4}\]

OpenStudy (anonymous):

The answer is 6 I agree with @GIL.ojei

OpenStudy (anonymous):

\[9-(\frac{ 1 }{ 16 })\times48\]

OpenStudy (anonymous):

48/16=3

OpenStudy (anonymous):

9-3=6